OpenStudy (anonymous):

integrate by parts: ∫√xln√x dx I know u=√x and dv=ln√x dx, I just need help with the rest!

OpenStudy (anonymous):

you should choose a better dv \[u = \ln{\sqrt{x}}\rightarrow du = \frac{dx}{2x}\] \[dv = \sqrt{x}dx\rightarrow v = \frac{2}{3}x^{\frac{3}{2}}\]

OpenStudy (anonymous):

will i find the same answer with the ones I chose? I worked it out while I was waiting and got -> (√x/x)-(x/√x)-(√x³/3)